# Bridging Graph and State-Space Modeling for Intensive Care Unit Length of Stay Prediction

### v1

@@ Abstract
Predicting a patient’s length of stay (LOS) in the intensive care unit (ICU) is a critical task for hospital resource management, yet remains challenging due to the heterogeneous and irregularly sampled nature of electronic health records (EHRs). In this work, we propose S2G-Net, a novel neural architecture that unifies state-space sequence modeling with multi-view Graph Neural Networks (GNNs) for ICU LOS prediction. The temporal path employs Mamba state-space models (SSMs) to capture patient trajectories, while the graph path leverages an optimized GraphGPS backbone, designed to integrate heterogeneous patient similarity graphs derived from diagnostic, administrative, and semantic features. Experiments on the large-scale MIMIC-IV cohort dataset show that S2G-Net consistently outperforms sequence models (BiLSTM, Mamba, Transformer), graph models (classic GNNs, GraphGPS), and hybrid approaches across all primary metrics. Extensive ablation studies and interpretability analyzes highlight the complementary contributions of each component of our architecture and underscore the importance of principled graph construction. These results demonstrate that S2G-Net provides an effective and scalable solution for ICU LOS prediction with multi-modal clinical data.
@@ 1 Introduction
There has been a growing focus on predicting patient outcomes in intensive care units (ICUs), along with increasing research interest in applying Graph Deep Learning techniques for this purpose Rocheteau et al. (2021). This surge in attention is largely driven by the expanded use of Electronic Health Records (EHRs), the demonstrated effectiveness of graph representation learning across diverse domains—including recommendation systems and misinformation detection Monti et al. (2019), financial forecasting Zhang et al. (2023), and gene regulatory network modeling de Ocáriz Borde et al. (2025), to name a few—and the broader effort to reduce preventable deaths through early warning systems Moon et al. (2011).
In this work, we bridge recent techniques in the Graph Neural Network (GNN) literature with sequence modeling methods by using graphs to model inter-patient correlations and sequence models to capture time-series patient data, enabling accurate predictions of ICU length of stay (LOS) for improved hospital resource allocation, management, and optimization. Despite recent progress, which we later discuss in Section 2.2, several challenges remain. Firstly, there is a clear gap in the literature for general-purpose models that jointly capture both temporal patient trajectories and inter-patient correlations, particularly for tasks like ICU LOS prediction, where existing approaches tend to focus on either temporal or spatial aspects in isolation or are tailored to narrow clinical outcomes such as mortality or disease progression. Secondly, traditional GNNs generally operate on single-view, static graphs, which are ill-suited for modeling heterogeneous, multi-modal clinical features and dynamically evolving patient relationships. Transformer-based graph backbones (e.g., GraphGPS Rampásek et al. (2022) and other Graph Transformers Sáez de Ocáriz Borde (2024)) are powerful but computationally demanding, with quadratic complexity due to all-to-all node-wise communication, and they often fail to accommodate the unique characteristics of ICU data, such as sparsity and multi-view structure. Moreover, many existing models lack sufficient interpretability Molnar et al. (2020) (which is also critical in many real-world ML applications such as engineering de Ocáriz Borde et al. (2022) or finance Brigo et al. (2021)), limiting their clinical relevance.
In this study, we introduce S2G-Net (State-space × Optimized Multi-View GNN) for ICU LOS prediction. It comprises three components: a temporal pathway that uses a state-space backbone (Mamba Gu and Dao (2023)) to handle irregular and long-range temporal dependencies in patient time-series data; a graph pathway that employs an optimized GraphGPS to learn from multiple patient similarity graphs constructed from diagnostic, semantic, and administrative views; and an auxiliary static-feature branch.
@@
  • We adapt and enhance the GraphGPS architecture for clinical use, replacing Transformer Vaswani et al. (2017) layers with state-space models (SSMs) and introducing more efficient graph fusion strategies.
  • We benchmark S2G-Net against strong sequence, graph, and hybrid baselines on the large-scale MIMIC-IV dataset Johnson et al. (2023), demonstrating superior performance in terms of accuracy, resource efficiency, and calibration.
@@ 2.1 Understanding the Complexities of Length of Stay Prediction
An ICU is a specialized healthcare unit providing advanced monitoring, physiologic organ support, and expert medical and nursing care for critically ill patients Marshall et al. (2017). Globally, ICU resources are limited and expensive. Bed capacity varies by region, from 4.2 beds per 100,000 people in Portugal to 29.2 in Germany Rhodes et al. (2012), with pandemics like COVID-19 pushing occupancy above 95% in some regions Grasselli et al. (2020). The daily cost per ICU patient cared for in developed countries ranges from \$2,000 to \$5,000 per day Chacko et al. (2023); Guest et al. (2020); Carrandi et al. (2024); Kılıç et al. (2019), while prolonged hospital stays significantly increase the financial burden on both hospitals and families. In the United States itself, ICU costs contribute nearly 20 per cent of hospital costs Wagner et al. (1983), amounting to over \$100 billion annually Kannan et al. (2023). Beyond the costs, ICU LOS is closely tied to patient outcomes. Prolonged LOS is related to complications like hospital acquired infections Toh et al. (2017), delirium, and muscle atrophy, while early discharge can lead to readmission or worsening of the prognosis Loreto et al. (2020). Research has shown that in developed healthcare systems, ICU stays typically range from 3 to 8 days Moitra et al. (2016), but vary widely depending on the condition. For example, patients recovering from cardiac surgery may require a 1-2 day hospital stay, whereas patients with sepsis or acute respiratory distress syndrome (ARDS) may require weeks of intensive care. Accurate LOS prediction can help reduce inefficiencies, improve patient care, and alleviate the financial and operational burden on healthcare systems.
@@
In fact, the dynamic and heterogeneous properties of ICU data make predicting ICU LOS inherently challenging. ICU patients are continuously monitored, generating high-frequency time-series data including vital signs (e.g., heart rate, blood pressure) Berry and Spain (2024), laboratory results (e.g., blood gases, inflammatory markers), and interventions (e.g., medications, mechanical ventilation). These data display complex and intricate dependencies over time and across organ systems. For example, oxygenation may improve with diuretics in patients with heart failure, whereas worsening renal function may coincide with hypotension or sepsis progression, creating a highly non-linear and time-dependent problem. Moreover, in real ICU settings, only partial data points are recorded at uniform intervals, with physiological parameters like continuously monitored heart rate and oxygen saturation providing high-frequency data, while laboratory test results like blood gas levels or electrolyte measurements are recorded intermittently. From a deep learning perspective, these characteristics present several challenges: ICU records are often irregular and incomplete, with certain variables sampled at widely varying frequencies. The data are also highly heterogeneous, varying across patients, conditions, and treatments. This makes it difficult for models to learn consistent patterns. Moreover, the relationships between variables are often non-linear and evolve over time, requiring models that can capture both short- and long-term dependencies. These challenges are further compounded by the need for models that are not only accurate but also interpretable and robust in real-world clinical settings.
@@ 2.2 Deep Learning Approaches for Clinical Data
Recurrent neural network (RNN) Rumelhart et al. (1986), particularly Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) models, are widely used for patient time-series modeling Liu et al. (2022); Ma et al. (2017); Rajkomar et al. (2018). For example, Harutyunyan et al. Harutyunyan et al. (2019) provided LSTM baselines for mortality, decompensation, length of stay, and phenotype prediction on MIMIC-III Johnson et al. (2016). LSTMs can model sequential dependencies and have been adapted for multivariate and irregular EHR data using masking and imputation Lipton et al. (2015). However, their performance declines with long and highly irregular ICU sequences. Recently, transformer architectures have shown promise in EHR modeling by capturing long-range dependencies through attention mechanisms. Models such as BEHRT Li et al. (2020) and Med-BERT Rasmy et al. (2021) adapt transformers to structured clinical data, while recent approaches introduce positional encodings and sparse attention for handling irregular sampling Song et al. (2018). Nonetheless, transformers are computationally intensive due to quadratic self-attention and can overfit with limited clinical data. Structured SSMs, such as S4 and Mamba, complement RNNs and transformers. S4 Gu et al. (2021) achieves linear-time sequence modeling and maintains long-term dependencies using structured kernels. Mamba Gu and Dao (2023) introduces input-dependent recurrence for greater stability and efficiency on long sequences. Both have been applied to domains such as genomics and time-series analysis, but applications in high error rate ICU data remain rare.
@@
Patient similarity graphs can model complex multi-modal relationships between individuals. Parisot et al. Parisot et al. (2018) applied graph convolutional networks (GCNs) to patient graphs for disease classification, constructing edges from demographic and imaging features. Subsequent studies have extended graph-based methods Kipf (2016); de Ocáriz Borde et al. (2023a) to tasks such as readmission prediction Golmaei and Luo (2021), medication recommendation Shang et al. (2019), sepsis risk forecasting Yin et al. (2025), as well as Alzheimer’s disease and mild cognitive impairment prediction de Ocáriz Borde et al. (2023b), demonstrating improvements over models that treat samples independently. Recent advances include dynamic and heterogeneous GNNs that capture temporal and attribute changes within patient cohorts. Dynamic GNNs model evolving edges and features over time Zheng et al. (2025), whereas multi-view GNNs integrate diverse data modalities (e.g., ICD codes, labs, text) for richer representations Kim et al. (2023); Liu et al. (2020). Nevertheless, most clinical applications to date still use static or single-view graphs due to computational and modeling complexity. Beyond these considerations, another challenge lies in balancing local neighborhood aggregation and global context modeling. Local-Global GNNs such as GraphGPS Rampásek et al. (2022) and GLGNN Eliasof and Treister (2024) address this issue by combining message passing with global attention modules. These approaches are well-suited to clinical graphs, where both local similarity and global cohort structure are relevant for prediction.
@@
Models that jointly learn from temporal patient trajectories and inter-patient graph structure have shown improved performance in clinical forecasting tasks. GAMENet Shang et al. (2019) integrates GCNs with dynamic memory modules for medication recommendation, and LSTM-GNN architectures Rocheteau et al. (2021) combine sequence modeling with graph-based neighborhood aggregation for ICU outcome prediction. Despite progress, most existing models inadequately integrate temporal and relational information, lack multi-modal graph construction, or are computationally inefficient. We address these gaps with S2G-Net, a unified and efficient hybrid framework.
@@ 3.2 Patient Graph Construction
We encode structured clinical similarity by constructing a graph from filtered ICD-9/10 codes. Each patient is represented by a sparse binary vector $\mathbf{d}_{s}\in\{0,1\}^{d_{d}}$, indicating assigned specific diagnosis. Edges are constructed based on pairwise similarity using one of three strategies: (i) TF-IDF Cosine Similarity, where diagnosis vectors are reweighted by inverse document frequency and cosine similarity selects the top-$k$ neighbors; (ii) FAISS Approximate KNN Johnson et al. (2019), which performs efficient inner product search over large patient sets; and (iii) Penalized Co-occurrence, which favors patients with high diagnostic overlap while penalizing multi-morbidity noise. After similarity calculation, we prune each node to its top-$k_{\text{diag}}$ neighbors and normalize edge weights via $\log(1+x)$ or z-score. We remove the bottom 30% of weak connections to improve sparsity. The resulting graph is denoted $\mathcal{G}_{\text{diag}}=(\mathcal{V},\mathcal{E}_{\text{diag}})$.
@@
To capture semantic similarity beyond diagnosis code overlap, we construct a graph using sentence embeddings derived from patients’ diagnosis descriptions. For each patient, diagnosis codes are concatenated into a single text sequence, truncated or zero-padded to 512 tokens. A pre-trained DistilBERT model Sanh et al. (2019); Devlin et al. (2019) encodes each input into a fixed-length embedding $\mathbf{b}_{s}\in\mathbb{R}^{768}$. Patient similarity is computed via Gaussian-kernelized Euclidean distance:
@@
where $\sigma$ is set to the median pairwise distance: $\sigma=\text{median}_{i\neq j}(|\mathbf{b}_{i}-\mathbf{b}_{j}|_{2})/\sqrt{2}$, following Zelnik-Manor and Perona (2004). Top-$k_{\text{bert}}$ neighbors are retained, and the resulting graph is $\mathcal{G}_{\text{bert}}=(\mathcal{V},\mathcal{E}_{\text{bert}})$.
@@
The initial graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is formed by combining $\mathcal{E}_{\text{diag}}$ and $\mathcal{E}_{\text{bert}}$, preserving edge types to enable modality-aware message passing. To ensure global connectivity and facilitate modelling long-range dependencies Liang et al. (2025), we optionally apply two rewiring strategies: MST-based bridging to connect disconnected components, and GDC-based diffusion using first-order Personalized PageRank Gasteiger et al. (2019). A maximum out-degree of 15 is enforced via stratified sampling over edge types to preserve edge diversity.
@@ 4 Experimental Validation
We evaluate S2G-Net on MIMIC-IV v3.1 dataset Johnson et al. (2023), a large-scale collection of de-identified ICU records from Beth Israel Deaconess Medical Center. It comprises 65,347 adult patients. We extract 216 features, including 174 time-series and 42 static, aligned to hourly bins within a 48-hour window. Missing values are imputed via forward fill and decay masking. The dataset is split patient-wise into train (70%), validation (15%), and test (15%).
@@
We compare our model with a baseline model suite covering three categories. (i) Sequential models include RNN Rumelhart et al. (1986), BiLSTM Graves and Schmidhuber (2005), and Transformer Vaswani et al. (2017) encoders for temporal sequence modeling, along with Mamba Gu and Dao (2023), a recent state-space model optimized for efficient long-range dependency learning. (ii) Graph-based models operate on static patient similarity graphs and include Graph Attention Network (GAT) Velickovic et al. (2017), Message Passing Neural Network (MPNN) Gilmer et al. (2017), GraphSAGE Hamilton et al. (2017), and Optimized GraphGPS Rampásek et al. (2022), which augments local message passing with global mixing via Mamba layers. (iii) Spatio-temporal models include LSTM-GNN Rocheteau et al. (2021), where patient time-series are encoded via LSTM and aggregated over static graphs, and dynamic LSTM-GNNs Rocheteau et al. (2021), which update graph structure during training to reflect evolving patient states. Results from spatial-temporal models differ from Rocheteau et al. Rocheteau et al. (2021), attributable to differences in datasets (MIMIC-IV vs. eICU) and preprocessing; they serve here as consistent baselines.
@@
We use hyperparameter tuning with Optuna Akiba et al. (2019) over 75 trials, choosing the best configuration by validation $R^{2}$. The optimal configuration adopts a 2-layer Mamba (128-dim), and a fusion coefficient of $\lambda=0.5$. The patient similarity graph is constructed using FAISS with $k_{\text{diag}}=3$, $k_{\text{BERT}}=1$, and MST.

### v2

# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

@@ Abstract
Predicting a patient’s length of stay (LOS) in the intensive care unit (ICU) is a critical task for hospital resource management, yet remains challenging due to the heterogeneous and irregularly sampled nature of electronic health records (EHRs). In this work, we propose S2G-Net, a novel neural architecture that unifies state-space sequence modeling with multi-view Graph Neural Networks (GNNs) for ICU LOS prediction. The temporal path employs Mamba state-space models (SSMs) to capture patient trajectories, while the graph path leverages an optimized GraphGPS backbone, designed to integrate heterogeneous patient similarity graphs derived from diagnostic, administrative, and semantic features. Experiments on the large-scale MIMIC-IV cohort dataset show that S2G-Net consistently outperforms sequence models (BiLSTM, Mamba, Transformer), graph models (classic GNNs, GraphGPS), and hybrid approaches across all primary metrics. Extensive ablation studies and interpretability analyzes highlight the complementary contributions of each component of our architecture and underscore the importance of principled graph construction. These results demonstrate that S2G-Net provides an effective and scalable solution for ICU LOS prediction with multi-modal clinical data. The code can be found at https://github.com/ShuqiZi1/S2G-Net.
@@ 1 Introduction
There has been a growing focus on predicting patient outcomes in intensive care units (ICUs), along with increasing research interest in applying Graph Deep Learning techniques for this purpose [53]. This surge in attention is largely driven by the expanded use of Electronic Health Records (EHRs), the demonstrated effectiveness of graph representation learning across diverse domains—including recommendation systems and misinformation detection [44], financial forecasting [66], and gene regulatory network modeling [11], to name a few—and the broader effort to reduce preventable deaths through early warning systems [45].
In this work, we bridge recent techniques in the Graph Neural Network (GNN) literature with sequence modeling methods by using graphs to model inter-patient correlations and sequence models to capture time-series patient data, enabling accurate predictions of ICU length of stay (LOS) for improved hospital resource allocation, management, and optimization. Despite recent progress, which we later discuss in Section 2.2, several challenges remain. Firstly, there is a clear gap in the literature for general-purpose models that jointly capture both temporal patient trajectories and inter-patient correlations, particularly for tasks like ICU LOS prediction, where existing approaches tend to focus on either temporal or spatial aspects in isolation or are tailored to narrow clinical outcomes such as mortality or disease progression. Secondly, traditional GNNs generally operate on single-view, static graphs, which are ill-suited for modeling heterogeneous, multi-modal clinical features and dynamically evolving patient relationships. Transformer-based graph backbones (e.g., GraphGPS [50] and other Graph Transformers [58]) are powerful but computationally demanding, with quadratic complexity due to all-to-all node-wise communication, and they often fail to accommodate the unique characteristics of ICU data, such as sparsity and multi-view structure. Moreover, many existing models lack sufficient interpretability [43] (which is also critical in many real-world ML applications such as engineering [12] or finance [4]), limiting their clinical relevance.
In this study, we propose S2G-Net (State-space × Optimized Multi-View GNN) for ICU LOS prediction. It comprises three components: a temporal pathway using a state-space backbone (Mamba [20]) to handle irregular and long-range temporal dependencies in patient time-series data; a graph pathway employing an optimized GraphGPS to learn from multiple patient similarity graphs built from diagnostic, semantic, and administrative views; and an auxiliary static-feature branch.
@@
  • We adapt and enhance the GraphGPS architecture for clinical use, replacing Transformer [60] layers with state-space models (SSMs) and introducing more efficient graph fusion strategies.
  • We benchmark S2G-Net against strong sequence, graph, and hybrid baselines on the large-scale MIMIC-IV dataset [27], demonstrating superior performance in terms of accuracy, resource efficiency, and calibration.
@@ 2.1 Understanding the Complexities of Length of Stay Prediction
An ICU is a specialized healthcare unit providing advanced monitoring, physiologic organ support, and expert medical and nursing care for critically ill patients [41]. Globally, ICU resources are limited and expensive. Bed capacity varies by region, from 4.2 beds per 100,000 people in Portugal to 29.2 in Germany [52], with pandemics like COVID-19 pushing occupancy above 95% in some regions [18]. The daily cost per ICU patient cared for in developed countries ranges from \$2,000 to \$5,000 per day [6, 22, 5, 33], while prolonged hospital stays significantly increase the financial burden on both hospitals and families. In the United States itself, ICU costs contribute nearly 20 per cent of hospital costs [62], amounting to over \$100 billion annually [30]. Beyond the costs, ICU LOS is closely tied to patient outcomes. Prolonged LOS is related to complications like hospital acquired infections [59], delirium, and muscle atrophy, while early discharge can lead to readmission or worsening of the prognosis [39]. Research has shown that in developed healthcare systems, ICU stays typically range from 3 to 8 days [42], but vary widely depending on the condition. For example, patients recovering from cardiac surgery may require a 1-2 day hospital stay, whereas patients with sepsis or acute respiratory distress syndrome (ARDS) may require weeks of intensive care. Accurate LOS prediction can help reduce inefficiencies, improve patient care, and alleviate the financial and operational burden on healthcare systems.
@@
In fact, the dynamic and heterogeneous nature of ICU data makes predicting ICU LOS inherently challenging. ICU patients are continuously monitored, generating high-frequency time-series data including vital signs (e.g., heart rate, blood pressure) [3], laboratory results (e.g., blood gases, inflammatory markers), and interventions (e.g., medications, mechanical ventilation). These data display complex and intricate dependencies over time and across organ systems. For example, oxygenation may improve with diuretics in patients with heart failure, whereas worsening renal function may coincide with hypotension or sepsis progression, creating a highly non-linear and time-dependent problem. Moreover, in real ICU settings, only partial data points are recorded at uniform intervals, with physiological parameters like continuously monitored heart rate and oxygen saturation providing high-frequency data, while laboratory test results like blood gas levels or electrolyte measurements are recorded intermittently. From a deep learning perspective, these characteristics present several challenges: ICU records are often irregular and incomplete, with certain variables sampled at widely varying frequencies. The data are also highly heterogeneous, varying across patients, conditions, and treatments. This makes it difficult for models to learn consistent patterns. Moreover, the relationships between variables are often non-linear and evolve over time, requiring models to capture both short- and long-term dependencies. These challenges are further complicated by the need for models to be accurate, interpretable, and robust in real-world clinical settings.
@@ 2.2 Deep Learning Approaches for Clinical Data
Recurrent neural network (RNN) [54], particularly Long Short-Term Memory (LSTM) [26] models, are widely used for patient time-series modeling [37, 40, 49]. For example, Harutyunyan et al. [24] provided LSTM baselines for mortality, decompensation, length of stay, and phenotype prediction on MIMIC-III [28]. LSTMs can model sequential dependencies and have been adapted for multivariate and irregular EHR data using masking and imputation [36]. However, their performance declines with long and highly irregular ICU sequences. Recently, transformer architectures have shown promise in EHR modeling by capturing long-range dependencies through attention mechanisms. Models such as BEHRT [34] and Med-BERT [51] adapt transformers to structured clinical data, while recent approaches introduce positional encodings and sparse attention for handling irregular sampling [57]. Nonetheless, transformers are computationally intensive due to quadratic self-attention and can overfit with limited clinical data. Structured SSMs, such as S4 and Mamba, complement RNNs and transformers. S4 [21] achieves linear-time sequence modeling and maintains long-term dependencies using structured kernels. Mamba [20] introduces input-dependent recurrence for greater stability and efficiency on long sequences. Both have been applied to domains such as genomics and time-series analysis, but applications in high error rate ICU data remain rare.
@@
Patient similarity graphs can model complex multi-modal relationships between individuals. Parisot et al. [46] applied graph convolutional networks (GCNs) to patient graphs for disease classification, constructing edges from demographic and imaging features. Subsequent studies have extended graph-based methods [32, 9] to tasks such as readmission prediction [17], medication recommendation [56], sepsis risk forecasting [64], as well as Alzheimer’s disease and mild cognitive impairment prediction [10], demonstrating improvements over models that treat samples independently. Recent advances include dynamic and heterogeneous GNNs that capture temporal and attribute changes within patient cohorts. Dynamic GNNs model evolving edges and features over time [67], whereas multi-view GNNs integrate diverse data modalities (e.g., ICD codes, labs, text) for richer representations [31, 38]. Nevertheless, most clinical applications to date still use static or single-view graphs due to computational and modeling complexity. Beyond these considerations, another challenge lies in balancing local neighborhood aggregation and global context modeling. Local-Global GNNs such as GraphGPS [50] and GLGNN [14] address this issue by combining message passing with global attention modules. These approaches are well-suited to clinical graphs, where both local similarity and global cohort structure are relevant for prediction.
@@
Models that jointly learn from temporal patient trajectories and inter-patient graph structure have shown improved performance in clinical forecasting tasks. GAMENet [56] integrates GCNs with dynamic memory modules for medication recommendation, and LSTM-GNN architectures [53] combine sequence modeling with graph-based neighborhood aggregation for ICU outcome prediction. Despite progress, most existing models inadequately integrate temporal and relational information, lack multi-modal graph construction, or are computationally inefficient. We address these gaps with S2G-Net, a unified and efficient hybrid framework.
@@ 3.2 Patient Graph Construction
We encode structured clinical similarity by constructing a graph from filtered ICD-9/10 codes. Each patient is represented by a sparse binary vector $\mathbf{d}_{s}\in\{0,1\}^{d_{d}}$, indicating assigned specific diagnosis. Edges are constructed based on pairwise similarity using one of three strategies: (i) TF-IDF Cosine Similarity, where diagnosis vectors are reweighted by inverse document frequency and cosine similarity selects the top-$k$ neighbors; (ii) FAISS Approximate KNN [29], which performs efficient inner product search over large patient sets; and (iii) Penalized Co-occurrence, which favors patients with high diagnostic overlap while penalizing multi-morbidity noise. After similarity calculation, we prune each node to its top-$k_{\text{diag}}$ neighbors and normalize edge weights via $\log(1+x)$ or z-score. We remove the bottom 30% of weak connections to improve sparsity. The resulting graph is denoted $\mathcal{G}_{\text{diag}}=(\mathcal{V},\mathcal{E}_{\text{diag}})$.
@@
To capture semantic similarity beyond diagnosis code overlap, we construct a graph using sentence embeddings derived from patients’ diagnosis descriptions. For each patient, diagnosis codes are concatenated into a single text sequence, truncated or zero-padded to 512 tokens. A pre-trained DistilBERT model [55, 13] encodes each input into a fixed-length embedding $\mathbf{b}_{s}\in\mathbb{R}^{768}$. Patient similarity is computed via Gaussian-kernelized Euclidean distance:
@@
where $\sigma$ is set to the median pairwise distance: $\sigma=\text{median}_{i\neq j}(|\mathbf{b}_{i}-\mathbf{b}_{j}|_{2})/\sqrt{2}$, following [65]. Top-$k_{\text{bert}}$ neighbors are retained, and the resulting graph is $\mathcal{G}_{\text{bert}}=(\mathcal{V},\mathcal{E}_{\text{bert}})$.
@@
The initial graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is formed by merging $\mathcal{E}_{\text{diag}}$ and $\mathcal{E}_{\text{bert}}$, retaining edge types for modality-aware message passing. To ensure global connectivity and capture long-range dependencies [35], we optionally apply two rewiring strategies: MST-based bridging to link disconnected components, and GDC-based diffusion using first-order Personalized PageRank [15]. A maximum out-degree of 15 is enforced via stratified edge-type sampling to preserve diversity.
@@ 4 Experimental Validation
We evaluate S2G-Net on MIMIC-IV v3.1 dataset [27], a large-scale collection of de-identified ICU records from Beth Israel Deaconess Medical Center. It comprises 65,347 adult patients. We extract 216 features, including 174 time-series and 42 static, aligned to hourly bins within a 48-hour window. Missing values are imputed via forward fill and decay masking. The dataset is split patient-wise into train (70%), validation (15%), and test (15%).
@@
We compare our model with a baseline model suite covering three categories. (i) Sequential models include RNN [54], BiLSTM [19], and Transformer [60] for temporal sequence modeling, along with Mamba [20], a recent state-space model optimized for efficient long-range dependency learning. (ii) Graph-based models operate on static patient similarity graphs, including Graph Attention Network (GAT) [61], Message Passing Neural Network (MPNN) [16], GraphSAGE [23], and Optimized GraphGPS [50], which augments local message passing with global mixing via Mamba layers. (iii) Spatio-temporal models include LSTM-GNN [53], where patient time-series are encoded via LSTM and aggregated over static graphs, and dynamic LSTM-GNNs [53], which update graph structure during training to reflect evolving patient states. Results differ from Rocheteau et al. due to dataset (MIMIC-IV vs. eICU) and preprocessing differences but are used here as consistent baselines. For completeness, XGBoost [7] is included as a non-neural baseline for structured clinical prediction.
@@
We tune hyperparameters with Optuna [1] over 75 trials, choosing the best configuration by validation $R^{2}$. The optimal setup adopts a 2-layer Mamba (128-dim) with a fusion coefficient of $\lambda=0.5$. The patient similarity graph is built via FAISS using $k_{\text{diag}}=3$, $k_{\text{BERT}}=1$, and MST.